OpenStudy (anonymous):

i am looking to figure out ratios in geometry. If the angle of a trianle are in the extended ratio 1:3:5, find their measures

OpenStudy (saifoo.khan):

x+3x+5 = 180.

OpenStudy (saifoo.khan):

Or, \[\frac{x}{8} \times 180\]

OpenStudy (saifoo.khan):

where x is the ratio.

OpenStudy (anonymous):

ok - let me try that - give me a moment

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

why isn't it 5x

OpenStudy (saifoo.khan):

typo.. sorry. that's 5x/

OpenStudy (anonymous):

so it should be set up as x + 3x + 5x = 180

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

so then the next step is 4x + 5x = 180 and then it would be 9x = 180 and you divide it from there

OpenStudy (saifoo.khan):

Correct. u will get. x = 20. so the smallest angle is 20.

OpenStudy (anonymous):

so then you subtract that from 180 and get 160 for the other angles

OpenStudy (saifoo.khan):

No. now. u know the value of two other angles. that is 3x and 5x. and you know the value of x for well. so, 3(x) = 3(20) = ? degrees/ And 5(x) = 5(20) = ? degress

OpenStudy (anonymous):

so then i would get 60, 100, and 20.

OpenStudy (saifoo.khan):

Wow, you got them right. :D

OpenStudy (anonymous):

:) thank you so much for ur help.

OpenStudy (saifoo.khan):

You welcome.

OpenStudy (anonymous):

do u have time to help me with a quadrilateral?

OpenStudy (saifoo.khan):

idk. i will see. if i can solve them for sure.

OpenStudy (anonymous):

alright. it has one angle of x+20, another of x-50, third one x+2, the last is x

OpenStudy (saifoo.khan):

post that on left please.

OpenStudy (anonymous):

ok. thanks again.